# Origin of the hidden energy scale and the $f$-ratio in geometrically frustrated magnets

### v3

@@ Abstract
Figure 1: Heat capacity $C(T)$ of the spin-1/2 $XXZ$ model on a cluster of 18 sites on the kagome lattice computed from exact diagonalization with periodic boundary conditions for various values of the anisotropy parameter $\alpha$ [cf. Eq. (3a)]. The left inset shows $C(T)$ at very low temperatures, which may display a secondary low-temperature peak. The right inset shows the $\alpha$-dependence of the primary ($T^{*}$) and secondary ($T_{L}$) low-temperature peak positions. The existence of distinguishable low- and high-temperature peaks in $C(T)$ is in agreement with the results of Refs. [11, 12, 1, 51, 31, 47, 48, 57, 28, 32, 53, 30, 54] for the isotropic Heisenberg model ($\alpha=1$) and of Refs. [1, 54] for the $XXZ$ models. As we clarify in the Suppmental Material [4], the peaks may merge and turn into a plateau for $\alpha\sim 1$ and other boundary conditions.
Figure 2: Illustrations of Ising ground state hyridization via the transverse coupling ^$\hat{\cal H}_{\text{XY}}$ on the a) triangular, b) kagome, and c) pyrochlore lattices. Each picture shows two Ising ground states which are hybridized by ^$\hat{\cal H}_{\text{XY}}$. Fully-filled (open) circles indicate spin-up (down) on the corresponding sites, and are the same between the two states. Half-filled circles correspond to the upper spin orientations in one Ising state and the lower orientations in the other.
@@
^=^$\hat{\cal H}=\hat{\cal H}_{\text{Ising}}+\hat{\cal H}_{\text{XY}}+\hat{\cal H}_{\text{disorder}}+\hat{\cal H}_{\text{weak}}$ (3a)
^$\hat{\cal H}_{\text{Ising}}=J\sum_{(ij)}{\hat{S}}_{i}^{z}{\hat{S}}_{j}^{z}$ (3b)
^$\hat{\cal H}_{\text{XY}}=\alpha J\sum_{(ij)}\left({\hat{S}}_{i}^{x}{\hat{S}}_{j}^{x}+{\hat{S}}_{i}^{y}{\hat{S}}_{j}^{y}\right)$ (3c)
^$\hat{\mathcal{H}}_{\mathrm{disorder}}=\sum_{(ij)}\delta J_{ij}{\hat{S}}_{i}^{z}{\hat{S}}_{j}^{z}$ (3d)
of spins on a GF lattice, where the part ^$\hat{\cal H}_{\text{Ising}}+\hat{\cal H}_{\text{XY}}$ describes a Heisenberg $XXZ$ model; the summation is carried out over all pairs $(ij)$ of neighbouring spins; $J$ is the coupling strength between the spins’ $z$ components; ^$\hat{\cal H}_{\text{disorder}}$ is a perturbation created by weak quenched disorder whose exact form is not important but is considered, for concreteness, to come from the random fluctuations $\delta J_{ij}$ of the exchange couplings; ^$\hat{\cal H}_{\text{weak}}$ accounts for weak non-exchange, e.g. dipole-dipole, interactions with characteristic energies significantly smaller than $J$, and the anisotropy parameter $\alpha$ describes the relative strength of the transverse and longitudinal exchange interactions.
At $\alpha=0$ (the Ising limit) and ^$\hat{\cal H}_{\text{disorder}}=\hat{\cal H}_{\text{weak}}=0$, spins on a GF lattice have extensive degeneracy, i.e. degeneracy that scales exponentially with the system size, whereas single spin-flip excitations have a characteristic energy of the order of $\theta_{W}\sim JZ$.
@@
Indeed, the hybridisation of the ground states of the Ising model at non-zero $\alpha$ is caused by the transverse-coupling Hamiltonian ^$\hat{\cal H}_{\text{XY}}$ given by Eq. (3c). The $n$-th order transition amplitude between two Ising ground states |$\ket{G_{1}}$ and |$\ket{G_{2}}$ can be estimated as
@@
The characteristic eigenvalues of the operator $\hat{\tilde{H}}_{\text{Ising}}$ in Eq. (4) are given by $\theta_{W}\sim ZJ$, while the matrix elements of the operators ^$\hat{\cal H}_{\text{XY}}$ are given by $\alpha J/2$. The transition amplitude between the Ising ground states due to $Z_{1}$ pairwise spin-exchange processes can, therefore, be estimated as
@@
The quantity (5) describes the characteristic energy splitting between the lowest energy levels of the $XXZ$ model with the Hamiltonian (3a)-(3c) in the limit of weak quenched disorder and weak non-exchange interactions. These states are adiabatically connected to the Ising ground states on the same lattice when the parameter $\alpha$ increases from 00 to a nonzero value. The quantity $Z_{1}$ in Eq. (5) is given by the minimum number of pairwise spin-exchange processes that can connect two Ising ground states. The value of $Z_{1}$, like $Z$, depends on the details of the lattice.
@@
For common $2$D lattices, the values of the respective Ising ground-state entropies can be obtained analytically, while for Ising models on other lattices, the entropy values are known from numerical simulations, as summarised in Table 2. Because the values of those (Ising) entropies match the entropies corresponding to the lower peaks in the more complicated, realistic model with the Hamiltonian (3a)-(3d), the value of that entropy of either peak of $C(T)$ can be used as an experimental test of the origin of the hidden energy scale.
@@ I Boundary conditions and the heat capacity of spins on the kagome lattice
Figure S2: Heat capacity $C(T)$ (per spin) of the spin-$1/2$ $XXZ$ model on cluster (b) of Fig. S1. For large values of $\alpha$, there are not two well-separated peaks, in contrast to heat capacities on cluster (a) and in Refs. [1, 22].
@@ II Derivation of the critical disorder strength for glass freezing
^=^$\hat{\cal H}=\hat{\cal H}_{0}+\hat{\mathcal{H}}_{\mathrm{disorder}}$ (S1a)
^$\hat{\mathcal{H}}_{\mathrm{disorder}}=\sum_{(ij)}\delta J_{ij}{\hat{S}}_{i}^{z\gamma}{\hat{S}}_{j}^{z\gamma}$ (S1b)
where ^$\hat{\cal H}_{0}$ is the Hamiltonian of $R$ clean, independent replicas of the system; ^$\hat{\mathcal{H}}_{\mathrm{disorder}}$ describes quenched fluctuations of the $zz$-couplings and $\gamma$ is the replica index. In the main text, the disorder-free Hamiltonian of the system is given by ^$\hat{\cal H}_{\mathrm{Ising}}+\hat{\cal H}_{\mathrm{XY}}$ [see Eqs. (3b) and (3c)] in a single replica; however, the exact form of ^$\hat{\cal H}_{0}$ is not important for the present calculation. We consider each replica to consist of a finite lattice of $N$ sites with periodic boundary conditions and coordination number $Z$. The sum in Eq. (S1b) runs over all nearest-neighbor site pairs $(ij)$. Hereinafter, summation over repeated indices is implied.
Using the interaction representation, where we take Eq. (S1b) as a perturbation to ^$\hat{\cal H}_{0}$, we can write the partition function of the $R$ replicas for a particular realization of the bond fluctuations in the form
@@
where $\beta$ is the inverse temperature, $Z_{0}$ is the partition function of a single clean replica, ${\hat{S}}_{i}^{z\gamma}(\tau)$ is the spin operator in the interaction representation, ${\hat{T}}_{\tau}$ is the Matsubara-time-ordering operator, and the angle brackets denote equilibrium thermal averaging with respect to ^$\hat{\cal H}_{0}$.
@@
where in writing the second line of Eq. (S3) we have introduced the matrix ${K}$ whose entries are $K_{ij}=1$ if sites $i$ and $j$ are first neighbors, and 00 otherwise.
@@
where the now explicit replica sums run over all pairs $(\gamma\delta)$ of replicas. The ${\cal O}(q)$ term in Eq. (S5) vanishes since clean replicas are independent of one another and the Hamiltonian ^$\hat{\cal H}_{0}$ is assumed to have global spin inversion symmetry. For the ${\cal O}(q^{2})$ term, the replica pair sums give two types of contributions: those in which either $\gamma_{2}=\gamma_{1}$, $\delta_{2}=\delta_{1}$ or $\gamma_{2}=\delta_{1}$, $\delta_{2}=\gamma_{1}$ and those in which there are three or four distinct replica indices. The latter type of term vanishes for the same reason as the ${\cal O}(q)$ contribution. The former type of term, however, does not vanish, and yields
@@
$-\beta{\cal F}(q)=\underset{R\rightarrow 0}{\mathrm{lim}}\frac{\overline{Z}^{R}-1}{R}=\frac{1}{4}\varkappa\beta^{2}q^{2}ZN\biggl{[}\frac{1}{2}-\varkappa Z\sum_{\ell}\int_{0}^{\beta}\mathrm{d}\tau\,\mathrm{d}\tau^{\prime}\,g(0,0;\ell,\tau)g(0,0;\ell,\tau^{\prime})\biggr{]}+{\cal O}(q^{3})$ (S9)
@@ II.1 Critical disorder strength in the effective Ising regime
where the subscript 00 on the state indices denotes a state which is adiabatically connected to an Ising ground state. We can decompose Eq. (S13) further as follows: writing
@@
where $\hat{\mathcal{P}}_{E}$ projects onto the excitations of ^$\hat{\cal H}_{0}$ which are adiabatically connected to the excited states of the Ising model, Eq. (S13) becomes
@@ II.2 Critical disorder strength at very low temperatures
assuming the Hamiltonian ^$\hat{\cal H}_{0}$ has a non-degenerate ground state.
@@
similarly to Bloch’s states in the single-particle case, where $|u\rangle$ is a state invariant under simultaneous translation of all lattice sites by a lattice vector. Hence, defining ${\hat{T}}_{\ell}$ as the translation operator which translates states by the lattice vector ${\bf R}_{\ell}$ connecting sites $\ell$ and 00, we have
@@
If the excitations remain gapped in the thermodynamic limit, the integral converges, and $\varkappa_{c}$ remains finite as $T\rightarrow 0$. If, however, the excitations are gapless, the integral will converge for $\rho(\xi)$ scaling faster than linearly at small $\xi$, and diverge otherwise. This means that in a gapless $3$D system with the power-law excitation dispersion $\xi(k)\propto k^{p}$ in the long-wave limit $k\rightarrow 0$,

### v4

@@ Abstract
Figure 1: Heat capacity $C(T)$ of the spin-1/2 XXZXXZ model on a cluster of 18 sites on the kagome lattice computed from exact diagonalization with periodic boundary conditions for various values of the anisotropy parameter $\alpha$ [cf. Eq. (3a)]. The left inset shows $C(T)$ at very low temperatures, which may display a secondary low-temperature peak. The right inset shows the $\alpha$-dependence of the primary ($T^{*}$) and secondary ($T_{L}$) low-temperature peak positions. The existence of distinguishable low- and high-temperature peaks in $C(T)$ is in agreement with the results of Refs. [11, 12, 1, 54, 32, 48, 50, 60, 29, 33, 56, 31, 57] for the isotropic Heisenberg model ($\alpha=1$) and of Refs. [1, 57] for the XXZXXZ models. As we clarify in the Suppmental Material [4], the peaks may merge and turn into a plateau for $\alpha\sim 1$ and other boundary conditions.
Figure 2: Illustrations of Ising ground state hyridization via the transverse coupling $\hat{\cal H}_{\text{XY}}$ on the a) triangular, b) kagome, and c) pyrochlore lattices. Each picture shows two Ising ground states which are hybridized by $\hat{\cal H}_{\text{XY}}$. Fully-filled (open) circles indicate spin-up (down) on the corresponding sites, and are the same between the two states. Half-filled circles correspond to the upper spin orientations in one Ising state and the lower orientations in the other.
@@
$\hat{\cal H}=\hat{\cal H}_{\text{Ising}}+\hat{\cal H}_{\text{XY}}+\hat{\cal H}_{\text{disorder}}+\hat{\cal H}_{\text{weak}}$ (3a)
$\hat{\cal H}_{\text{Ising}}=J\sum_{(ij)}{\hat{S}}_{i}^{z}{\hat{S}}_{j}^{z}$ (3b)
$\hat{\cal H}_{\text{XY}}=\alpha J\sum_{(ij)}\left({\hat{S}}_{i}^{x}{\hat{S}}_{j}^{x}+{\hat{S}}_{i}^{y}{\hat{S}}_{j}^{y}\right)$ (3c)
$\hat{\mathcal{H}}_{\mathrm{disorder}}=\sum_{(ij)}\delta J_{ij}{\hat{S}}_{i}^{z}{\hat{S}}_{j}^{z}$ (3d)
of spins on a GF lattice, where the part $\hat{\cal H}_{\text{Ising}}+\hat{\cal H}_{\text{XY}}$ describes a Heisenberg $XXZ$ model; the summation is carried out over all pairs $(ij)$ of neighbouring spins; $J$ is the coupling strength between the spins’ $z$ components; $\hat{\cal H}_{\text{disorder}}$ is a perturbation created by weak quenched disorder whose exact form is not important but is considered, for concreteness, to come from the random fluctuations $\delta J_{ij}$ of the exchange couplings; $\hat{\cal H}_{\text{weak}}$ accounts for weak non-exchange, e.g. dipole-dipole, interactions with characteristic energies significantly smaller than $J$, and the anisotropy parameter $\alpha$ describes the relative strength of the transverse and longitudinal exchange interactions.
At $\alpha=0$ (the Ising limit) and $\hat{\cal H}_{\text{disorder}}=\hat{\cal H}_{\text{weak}}=0$, spins on a GF lattice have extensive degeneracy, i.e. degeneracy that scales exponentially with the system size, whereas single spin-flip excitations have a characteristic energy of the order of $\theta_{W}\sim JZ$.
@@
[…]
@@
The characteristic eigenvalues of the operator $\hat{\tilde{H}}_{\text{Ising}}$ in Eq. (4) are given by $\theta_{W}\sim ZJ$, while the matrix elements of the operators $\hat{\cal H}_{\text{XY}}$ are given by $\alpha J/2$. The transition amplitude between the Ising ground states due to $Z_{1}$ pairwise spin-exchange processes can, therefore, be estimated as
@@
The quantity (5) describes the characteristic energy splitting between the lowest energy levels of the $XXZ$ model with the Hamiltonian (3a)-(3c) in the limit of weak quenched disorder and weak non-exchange interactions. These states are adiabatically connected to the Ising ground states on the same lattice when the parameter $\alpha$ increases from 0 to a nonzero value. The quantity $Z_{1}$ in Eq. (5) is given by the minimum number of pairwise spin-exchange processes that can connect two Ising ground states. The value of $Z_{1}$, like $Z$, depends on the details of the lattice.
@@
For common 22D lattices, the values of the respective Ising ground-state entropies can be obtained analytically, while for Ising models on other lattices, the entropy values are known from numerical simulations, as summarised in Table 2. Because the values of those (Ising) entropies match the entropies corresponding to the lower peaks in the more complicated, realistic model with the Hamiltonian (3a)-(3d), the value of that entropy of either peak of $C(T)$ can be used as an experimental test of the origin of the hidden energy scale.
@@ I Boundary conditions and the heat capacity of spins on the kagome lattice
Figure S2: Heat capacity $C(T)$ (per spin) of the spin-$1/2$ XXZXXZ model on cluster (b) of Fig. S1. For large values of $\alpha$, there are not two well-separated peaks, in contrast to heat capacities on cluster (a) and in Refs. [1, 22].
@@ II Derivation of the critical disorder strength for glass freezing
$\hat{\cal H}=\hat{\cal H}_{0}+\hat{\mathcal{H}}_{\mathrm{disorder}}$ (S1a)
$\hat{\mathcal{H}}_{\mathrm{disorder}}=\sum_{(ij)}\delta J_{ij}{\hat{S}}_{i}^{z\gamma}{\hat{S}}_{j}^{z\gamma}$ (S1b)
where $\hat{\cal H}_{0}$ is the Hamiltonian of $R$ clean, independent replicas of the system; $\hat{\mathcal{H}}_{\mathrm{disorder}}$ describes quenched fluctuations of the $zz$-couplings and $\gamma$ is the replica index. In the main text, the disorder-free Hamiltonian of the system is given by $\hat{\cal H}_{\mathrm{Ising}}+\hat{\cal H}_{\mathrm{XY}}$ [see Eqs. (3b) and (3c)] in a single replica; however, the exact form of $\hat{\cal H}_{0}$ is not important for the present calculation. We consider each replica to consist of a finite lattice of $N$ sites with periodic boundary conditions and coordination number $Z$. The sum in Eq. (S1b) runs over all nearest-neighbor site pairs $(ij)$. Hereinafter, summation over repeated indices is implied.
Using the interaction representation, where we take Eq. (S1b) as a perturbation to $\hat{\cal H}_{0}$, we can write the partition function of the $R$ replicas for a particular realization of the bond fluctuations in the form
@@
where $\beta$ is the inverse temperature, $Z_{0}$ is the partition function of a single clean replica, ${\hat{S}}_{i}^{z\gamma}(\tau)$ is the spin operator in the interaction representation, ${\hat{T}}_{\tau}$ is the Matsubara-time-ordering operator, and the angle brackets denote equilibrium thermal averaging with respect to $\hat{\cal H}_{0}$.
@@
where in writing the second line of Eq. (S3) we have introduced the matrix K{K} whose entries are $K_{ij}=1$ if sites $i$ and $j$ are first neighbors, and 0 otherwise.
@@
where the now explicit replica sums run over all pairs $(\gamma\delta)$ of replicas. The ${\cal O}(q)$ term in Eq. (S5) vanishes since clean replicas are independent of one another and the Hamiltonian $\hat{\cal H}_{0}$ is assumed to have global spin inversion symmetry. For the ${\cal O}(q^{2})$ term, the replica pair sums give two types of contributions: those in which either $\gamma_{2}=\gamma_{1}$, $\delta_{2}=\delta_{1}$ or $\gamma_{2}=\delta_{1}$, $\delta_{2}=\gamma_{1}$ and those in which there are three or four distinct replica indices. The latter type of term vanishes for the same reason as the ${\cal O}(q)$ contribution. The former type of term, however, does not vanish, and yields
@@
$\beta{\cal F}(q)=\underset{R\rightarrow 0}{\mathrm{lim}}\frac{\overline{Z}^{R}-1}{R}=\frac{1}{4}\varkappa\beta^{2}q^{2}ZN\biggl{[}\frac{1}{2}-\varkappa Z\sum_{\ell}\int_{0}^{\beta}\mathrm{d}\tau\,\mathrm{d}\tau^{\prime}\,g(0,0;\ell,\tau)g(0,0;\ell,\tau^{\prime})\biggr{]}+{\cal O}(q^{3})$ (S9)
@@ II.1 Critical disorder strength in the effective Ising regime
where the subscript 0 on the state indices denotes a state which is adiabatically connected to an Ising ground state. We can decompose Eq. (S13) further as follows: writing
@@
where $\hat{\mathcal{P}}_{E}$ projects onto the excitations of $\hat{\cal H}_{0}$ which are adiabatically connected to the excited states of the Ising model, Eq. (S13) becomes
@@ II.2 Critical disorder strength at very low temperatures
assuming the Hamiltonian $\hat{\cal H}_{0}$ has a non-degenerate ground state.
@@
similarly to Bloch’s states in the single-particle case, where $|u\rangle$ is a state invariant under simultaneous translation of all lattice sites by a lattice vector. Hence, defining $^{\hat{T}}_{\ell}$ as the translation operator which translates states by the lattice vector $\bf R$$_{\ell}$ connecting sites $\ell$ and 0, we have
@@
If the excitations remain gapped in the thermodynamic limit, the integral converges, and $\varkappa_{c}$ remains finite as $T\rightarrow 0$. If, however, the excitations are gapless, the integral will converge for $\rho(\xi)$ scaling faster than linearly at small $\xi$, and diverge otherwise. This means that in a gapless 33D system with the power-law excitation dispersion $\xi(k)\propto k^{p}$ in the long-wave limit $k\rightarrow 0$,